Tom is 11 years older than pat. The sum of their ages is 33. How old are they?
Answer by ikleyn(52786)

.

Tom   + Pat = 33


(P+11) + P  = 33


2P + 11 = 33


2P = 33 - 11 = 22


P = 22/2 = 11.


Answer.  Pat is 11 years old.  Tom is 22 years old. 

Solved.

Do you understand everything as it is presented in my solution ?